40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 3 22:41:17 2019, shruti, Update, General, PLL for PM measurement  PM_measurement.jpeg
    Reply  Wed Dec 4 20:32:43 2019, gautam, Update, General, PLL for PM measurement  
    Reply  Mon Dec 9 21:22:46 2019, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 
       Reply  Tue Dec 17 20:08:09 2019, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 20191217.png
          Reply  Thu Jan 16 19:32:23 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 
             Reply  Thu Jan 23 20:08:49 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT Image-1.jpg
                Reply  Tue Jan 28 19:40:15 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT PMTF.pdfBeatSpectrum.pdf
Message ID: 15148     Entry time: Thu Jan 23 20:08:49 2020     In reply to: 15129     Reply to this: 15169
Author: shruti 
Type: Update 
Category: General 
Subject: PLL / PM measurement of Xend NPRO PZT 

Setup Update:

- No more SR 560, upgraded to LB1005 P-I controller.  Because: Elog 14687. Schematic of new setup shown in Attachment 1.

- For this, the Marconi was moved to the other (east) side of the PSL table and a power supply was also placed in the enclosure.

I think that the RF power at the mixer in this new configuration is 0 dBm (since the spectrum analyzer read ~ -20 dBm)

Progress Today:

- Turned up the HEPA to 100%, closed the PSL shutter, misaligned the ITMX, connected the LB1005 to the PZT. [The PZT has been reconnected to the X arm PDH servo, HEPA back to 20-30%]

- Tried to look for the PSL+X beat, but it was not there. Gautam identified the flipmount in the path which sorted it out (eventually), but there was no elog about itsurprise.

- After much trial, the loop seemed to lock with PI corner 1 kHz, gain ~2.9 (as read on knob), LFGL set to 90 dB. The beat note looked quite stable on the oscilloscope, but the error signal had an rms of ~100 mV (Rana pointed out that it could be the laser noise) and the lock lasted for ~1 min each time.

The parameters were similar to that in elog 14687. Why do we require such a high PI corner frequency and LFGL?

Attachment 1: Image-1.jpg  189 kB  Uploaded Thu Jan 23 21:24:55 2020  | Hide | Hide all
Image-1.jpg
ELOG V3.1.3-